# Nonlinear Causal Discovery through a Sequential Edge Orientation Approach

### v1

@@ 3 Conduct variance test on $\widehat{\omega}_{n}^{2}$ (16) for model equivalence;
5       Compute the likelihood ratio test statistic $LR_{n}(\widehat{\theta},\widehat{\gamma})$ (15);
@@ 6       Obtain p-values $p_{X\rightarrow Y},p_{Y\rightarrow X}$ and preferred edge direction $E_{X,Y}:$
Our full algorithm involves constructing regression models in several tasks. This includes estimating residuals for computing $\widetilde{I}(X,Y)$ to rank edges, fitting models to compute the log-likelihood values under possible configurations in subgraphs, and performing covariate testing in the last stage. In the software implementation, the algorithm utilizes generalized additive models implemented in the mgcv package (Wood and Wood, 2015) to construct regression models, with the thin plate spline selected as the basis function.
@@ 5 Structural Learning Consistency
where $\mathcal{F}_{n}^{\oplus k},k=|S|$ is the space of additive functions defined in (6). We now list the relevant assumptions and formally state the consistency of the algorithm.
@@
Suppose $\{X_{i}\}_{i=1}^{p}$ follows the SEM (4) with a DAG $\mathcal{G}_{0}$ that satisfies the conditions for a restricted additive noise model and the following assumptions:
@@
The consistency of $\widehat{\mathcal{G}}_{n}$ also relies on the consistency of the statistical tests performed. Pertinent to skeleton learning in stage 1, Assumption 2 (A1) states there exists a lower bound $\tau>0$ for the partial correlation when $X_{i}\not\perp\!\!\!\!\perp X_{j}\mid S_{ij}$. We show that the probabilities of type I and type II errors converge to 0 for the CI tests in the large sample limit, by which our algorithm obtains a consistent estimate of the CPDAG. This assumption implies faithfulness between the DAG $\mathcal{G}_{0}$ and the joint probability distribution $p(X_{1},\ldots,X_{p})$ as well. Assumptions (A2) and (A3) are pertinent to identifying undirected edges that satisfy the PANM criterion. We assume the existence of a gap $\delta>0$ for $MI(\varepsilon_{i,S},X_{k}),k\in S$ to precisely distinguish edges that do and do not follow the PANM. Last, we assume a mild boundedness assumption on the entropy of each $X_{i}$ and various residual variables, which guarantees that the normalized independence measure (9) is well-defined.
@@
We analyze the computational complexity by counting the number of statistical tests performed and regression models fitted in the algorithm. For a $p-$node problem, the learned CPDAG can be a complete graph consisting of $p(p-1)/2$ edges in the worst-case scenario. The CPDAG $\mathcal{E}=(V,E)$ would generally contain much fewer edges and its undirected edges $\{U\}_{i=1}^{m},m<|E|$ would only account for a fraction of all edges. To compare two causal directions in the edge orientation procedure, our method builds two models for each direction and conducts one test per direction. The edge ranking procedure would perform $2m$ tests and fit $4m$ regression models, while the orientation procedure would at most perform $|U|=m$ tests and fit $4m$ models. However, there are fewer tests and models required in practice because Meek’s rules would orient additional edges. The computational complexity of procedure OrientEdges is then of order $O(m)$. In the edge deletion step, the method performs one significance test per node on its parent nodes, amounting to $p$ tests and $p$ models, and gives a complexity of order $O(p)$. While the PC algorithm only conducts conditional independence tests and is exponential with respect to $p$ in the worst case, it becomes polynomial when the underlying DAG is sparse. The empirical runtime comparisons are provided in Section 6.3.
6 Experiments
@@
We conducted numerical experiments to benchmark the accuracy and effectiveness of our method. At a detailed level, we assess the performance of the ranking procedure in Section 6.1, as well as the type I error rate and statistical power of the likelihood ratio test in Section 6.2. We then compare our method to competing causal discovery algorithms, using both simulated and real-world data sets in Sections 6.3 and 6.5. Intermediate results from each stage of our algorithm are provided in Section 6.4 to illustrate the effects of the individual components.
@@
We present two variations of the algorithm: the sample-splitting (SNOE-SS) approach, which is delineated in Algorithm 4, and the cross-validation (SNOE-CV) approach. The CV approach employs two-fold cross-validation in Algorithm 4 to perform the likelihood ratio test twice by exchanging the training and test data sets and uses either the smaller or larger p-value for evaluation. The larger p-value is used in our experiments, but practitioners may specify either option. To learn the initial graph, we implemented our modified version of the PC-stable algorithm (Colombo et al., 2014) from the bnlearn package coupled with the partial correlation test (Scutari, 2010). A more stringent threshold of $\alpha_{1}=0.05$ was applied for learning the CPDAG, while a relaxed threshold of $\alpha_{2}=0.25$ was used for obtaining the additional candidate edges. The significance level for the likelihood ratio test was set at $\alpha=0.05$ and further reduced to $\alpha=0.0001$ for edge pruning. The algorithm is implemented as an R package and can be accessed at https://github.com/stehuang/snoe.git.
@@ 6.1 Accuracy of Ranking Procedure
We performed several experiments to verify the precision of our edge ranking procedure. Specifically, we tested its ability to correctly rank undirected edges in a CPDAG. Three distinct DAG structures were considered, as depicted in Figure 4, with $N=200$ data sets containing $n=2000$ samples generated for each structure. As the data was generated using nonlinear functions, the edge directions can be recovered under all settings. Note that the cubic and piecewise functions are designed to be invertible. Given the CPDAG of each DAG, we computed the edge-wise independence measure $\widetilde{I}(X,Y)$, defined in (8), for each undirected edge $X-Y$ and ranked the edges in ascending order. We assessed whether an edge satisfying the PANM was ranked first.
$A$$B$$Z$$X$$Y$DAG 1
$A$$B$$Z$$X$$Y$DAG 2
$A$$B$$Z$$X$$Y$$W$$V$DAG 3
@@
Results are presented in Table 1 and are categorized by the data generating function used in the simulations. The values represent the proportion of data sets in which an undirected edge was ranked first for orientation. Uniformly across all graphs and functions, the edge following the PANM has the highest proportion of being ranked first. It is identified in the vast majority of datasets generated by the quadratic and sigmoid functions. Notably, our ranking procedure successfully determined the edge that satisfies the PANM in most cases under the cubic and piecewise functions, which is more challenging to distinguish since these invertible functions can be approximated by a linear function. These experiments verify both the ranking procedure and the use of the edge-wise independence measure to identify an edge fulfilling the orientation criterion.
@@
1 $Z-X$ Yes 0.64 0.62 0.99 0.72
1 $X-Y$ No 0.36 0.38 0.01 0.28
2 $Z-X$ Yes 0.43 0.52 0.46 0.93
2 $Z-Y$ No 0.34 0.25 0.35 0.07
2 $X-Y$ No 0.23 0.23 0.19 0
3 $Z-X$ Yes 0.38 0.52 0.95 0.41
3 $X-Y$ No 0.14 0.24 0.01 0.19
3 $Y-W$ No 0.26 0.14 0.01 0.21
3 $W-V$ No 0.22 0.11 0.03 0.19
Table 1: Breakdown of undirected edges ranked first under various settings.
@@ 6.2 Evaluation of the Likelihood Ratio Test
In this subsection, we investigate the type I error and statistical power of the likelihood ratio test. We considered five disinct DAG structures, each comprising of 2 to 5 nodes and 1 to 7 edges, and generated datasets of sample sizes $n=250,500,1000,1500,2000$. In the CPDAG of each network, we applied the likelihood ratio test to a targeted undirected edge to determine its causal direction. A total of $N=400$ tests were performed per graphical structure and sample size setting. Under a linear Gaussian DAG, the true edge direction of the targeted undirected edge is not identifiable. Under a nonlinear DAG, the true edge direction should be identifiable.
@@
The type I error of the test, averaged across all DAG structures per sample size, is documented in Figure 5. Overall, the type I error deviates minimally from the specified significance level and stabilizes as the sample size grows. Under both significance levels, the maximal difference between $\alpha$ and the type I error is within 0.015 for both the sample-splitting and cross-validation approaches. This also signifies that the test is robust against type I errors at smaller sample sizes, which can be attributed to the sample splitting design that renders the estimated model and log-likelihood estimates independent. The difference between the two approaches is minimal as well, differing by 0.005 at $n=2000$, and indicates that both effectively control the false positive rate.
@@
A type II error under the likelihood ratio test would be to falsely declare the two models as equivalent when only one model is true. In regards to edge orientation, that is to fail at identifying the true causal direction of an edge or incorrectly orient an edge. In this experiment, we applied the likelihood ratio test to data generated from a singular non-linear function, where the orientation of the targeted edge is identifiable. We recorded the power of the test under significance level $\alpha=0.05$.
As seen in Figure 6, the statistical power of the likelihood ratio test increases with the sample size across all function types. For certain functions like the piecewise and secant functions, the power approaches 1 with $n\geq 1500$ samples. For other functions, the power improves more gradually but still strictly increases with the sample size, including the case of the cubic function. In particular, the power increases by at least 33% between sample sizes $n=250$ and $n=1000$. The cross-validation based approach exhibits greater power, achieving at least 85% by $n=2000$ across all function types. The results provide empirical evidence that the test can in practice identify the true causal direction of an edge in a CPDAG, especially with sufficient data.
Figure 6: The statistical power was empirically measured by applying the test to an undirected edge in a CPDAG, with select nonlinear functions underlying the SEM. Results show that power increases as $n$ increases to at least 80%.
@@ 6.3 Comparison of Algorithm Performances
We evaluated the performance of SNOE against competing methods on synthetic data. We compared our method to CAM, NOTEARS, DAGMA, and SCORE, where each represents a different approach for learning nonlinear DAGs. CAM (Bühlmann et al., 2014) is a score-based method that assumes a nonlinear additive noise model (4) with Gaussian noise and optimizes the log-likelihood function to learn a DAG. Utilizing deep learning methods to model SEMs, NOTEARS (Zheng et al., 2020) and DAGMA (Bello et al., 2022) formulate the structural learning problem as a continuous-optimization problem with an algebraic constraint to enforce acyclicity. In this experiment, we employed the version tailored to learning from nonlinear data for both algorithms. SCORE (Rolland et al., 2022) employs a bottoms-up approach to iteratively identify leaf nodes by computing the Jacobian of the score function under the assumption of a Gaussian error distribution. For all methods, we used their recommended parameter settings. While several constraint-based methods were tested as well, their performance fell short. A detailed analysis of their performances is provided in Appendix C.1.
The algorithms were applied to learn six DAG structures of varying sizes selected from the bnlearn network repository. For each network, we generated $N=75$ data sets, each with a sample size of $n=1000$, using the additive noise model with Gaussian noise in Equation 4. The SEMs were created under three separate functional forms: the linear model, invertible functions, and non-invertible functions. The function classes are denoted respectively as linear, inv, and ninv in the figures. Under the invertible functions setting, we randomly selected functions from a set consisting of the cubic, inverse sine, piece-wise linear, and exponential functions. The non-invertible functions were sampled from a Gaussian process using a squared exponential kernel and bandwidth $h\sim\text{Unif}(5,5.25)$. For all cases, the Gaussian noise term was sampled with mean $\mu=0$ and standard deviation $\sigma\sim\text{Unif}(0.5,0.75)$.
The true DAG serves as the ground truth for evaluation, with the exception of the linear, Gaussian case for which only the equivalence class is identifiable and thus the true CPDAG is used as the ground truth. Our main evaluation metrics are the F1 score, structural Hamming distance (SHD), and computational complexity. The F1 score is a harmonic mean of the precision and recall scores. It is calculated as $F1=\frac{2TP}{2TP+FP+FN+IO}$, where TP, FP, FN, and IO respectively denote the number of true positives, false positives, false negatives, and incorrectly oriented edges. The SHD measures the number of edge additions, deletions, and reversals required to convert the learned DAG into the true DAG. The computational complexity is measured by overall runtime in seconds.
@@
The results presented in Figure 7 demonstrate that SNOE consistently outperforms competing methods. We observe that our algorithm achieved uniformly high F1 scores across all network structures and functional forms, with an average standard deviation of 0.06 in its performance across the three types of functions for both approaches. SNOE performed particularly well on invertible nonlinear DAGs, which presents a more challenging task due to the difficulty of detecting such nonlinear relations. On average across all function types, the F1 score of SNOE is respectively 67.6%, 61.8%, and 100.1% higher than those of NOTEARS, DAGMA, and SCORE. A closer analysis reveals that NOTEARS and DAGMA produced sparser DAGs by missing considerable amounts of edges, while SCORE often included many extraneous edges without capturing the true edges. CAM performed similarly to our method under data generated from non-invertible functions but showed considerable variability based on the data-generating function. Its F1 scores for invertible and linear functions were lower than SNOE with big margins for many cases. Specifically, CAM produced denser DAGs with more false positives, especially under these settings.
Our method also achieved significantly faster runtimes than the competing methods. As shown in Table 2, the average runtime and the standard deviation are reported respectively on the top and bottom rows in the cells. The sample-splitting approach (SNOE-SS) has the lowest runtime and smallest standard deviation, as expected, since the cross-validation approach performs regression twice. It is 8 times faster than all competing methods on the largest network (Magic). While CAM shows similar F1 scores in certain cases, SNOE-SS is 2.8 to 10.7 times faster than CAM overall. This efficiency can be attributed to its local learning approach of identifying edges satisfying the PANM, rather than optimizing over the entire DAG. Further reduction in runtime occurs as Meek’s rules typically orient additional undirected edges after performing the orientation test. In contrast to score-based and optimization-based methods that search over a restricted DAG space, our method leverages properties inherent to the graphical structure and only evaluates a sub-DAG, containing just the relevant nodes, to determine the correct edge orientations. Consequently, this results in higher computational efficiency for structural learning.
Algorithm Mehra Mildew Water Alarm Magic2 Magic CAM 95.29 (131.27) 189.42 (198.64) 228.93 (410.26) 138.93 (53.46) 215.68 (95.06) 895.07 (1082.33) DAGMA 741.15 (279.02) 822.97 (526.83) 1041.95 (2121.11) $1$.49 (986.63) 1448.05 (4347.48) 1834.14 (2339.3) NOTEARS 280.05 (479.05) 379.01 (541.79) 621.99 (1201.77) 846.99 (983.69) 592.57 (945.16) 1148.4 (1893.71) SCORE 40.55 (13.69) 239.1 (440.34) 96.95 (60.42) 198.89 (293.82) 243.18 (65.29) 639.46 (199.41) SNOE-SS 33.68 (14.24) 53.46 (54) 60.69 (76.01) 55.59 (51) 61.34 (56.92) 83.48 (58.81) SNOE-CV 37.99 (20.56) 55.86 (57.57) 63.65 (77.72) 60.56 (57.27) 68.09 (74.48) 95.79 (75.88)
Table 2: Average runtime of learning from data generated by all three function types in seconds on various networks; the standard deviation is reported in the parentheses.
@@
We further investigated the empirical performance of our algorithm under model mis-specification, specifically when the noise distribution is non-Gaussian, with results presented in Figure 8. Recall that the DAG is identifiable when the noise terms follow a non-Gaussian distribution. For this experiment, we simulated data under the same previous settings, but sampled the error variables from three different non-Gaussian distributions: the t-distribution with $df=5$, the Laplace distribution, and the Gumbel distribution, all with $\mu=0$ and $\sigma\sim\text{Unif}(0.5,0.75)$. Since the learning accuracies from each error distribution are similar, we present the combined results. Both variations of SNOE achieved higher accuracy than competing methods across all settings again, where only CAM is comparable in a few cases. Similar to the Gaussian case, the F1 score of our method is consistent under different function types with non-Gaussian noise. The exact F1 scores are comparable to the Gaussian case as well, indicating that our method is robust to model mis-specification and is a versatile causal learning method.

### v2

@@ 3 Conduct variance test on $\widehat{\omega}_{n}^{2}$ (16) for model equivalence;
[…]
@@ 6   Obtain p-values $p_{X\rightarrow Y},p_{Y\rightarrow X}$ and preferred edge direction $E_{X,Y}:$
While the final output of Algorithm 2 is a PDAG, users may specify to return a DAG if they believe that the nonlinear ANM assumption holds. Then, in an additional fourth and final stage, the edge orientation procedure in Algorithm 3 is applied again to extend the PDAG to a DAG. If any undirected edge still remains, the algorithm chooses the orientation with a higher log-likelihood value as the inferred causal direction of the edge.
Remark 3.
Our full algorithm involves constructing regression models in several tasks. This includes estimating residuals for computing $\widetilde{I}(X,Y)$ to rank edges, fitting models to compute the log-likelihood values under possible configurations in subgraphs, and performing covariate testing in the last stage. In the software implementation, the algorithm utilizes generalized additive models from in the mgcv package (mgcv_package) to construct regression models, with the thin plate spline selected as the basis function.
@@ 5 Structural Learning Consistency
where $k=|S|$ and $\mathcal{F}_{n}^{\oplus k}$ is the space of additive functions defined in (6). We now list the relevant assumptions and formally state the consistency of the algorithm.
@@
Suppose $\{X_{i}\}_{i=1}^{p}$ follows a restricted ANM (4) with a faithful DAG $\mathcal{G}_{0}$ and satisfies the following assumptions:
@@
The consistency of $\widehat{\mathcal{G}}_{n}$ also relies on the consistency of the statistical tests performed. Pertinent to skeleton learning in stage 1, Assumption 2 (A1) states that there exists a lower bound $\tau>0$ for the partial correlation when $X_{i}\not\perp\!\!\!\!\perp X_{j}\mid S_{ij}$. We show that the probabilities of type I and type II errors converge to 0 for the CI tests in the large sample limit, by which our algorithm obtains a consistent estimate of the CPDAG. Assumptions (A2) and (A3) are pertinent to identifying undirected edges that satisfy the PANM criterion. We assume the existence of a gap $\delta>0$ for $MI(\varepsilon_{i,S},X_{k}),k\in S$ to precisely distinguish edges that do and do not follow the PANM. Last, we assume a mild boundedness assumption on the entropy of each $X_{i}$ and various residual variables, which guarantees that the normalized independence measure (9) is well-defined.
@@
We analyze the computational complexity by counting the number of statistical tests performed and regression models fitted in the algorithm. For a $p$-node problem, the learned CPDAG can be a complete graph consisting of p(p1)/2p(p-1)/2 edges in the worst-case scenario. The CPDAG $\mathcal{E}=(V,E)$ generally contains much fewer edges and its undirected edges $\{U\}_{i=1}^{m},m<|E|$ generally account for only a fraction of all edges. To compare two causal directions in the edge orientation procedure, our method builds two models for each direction and conducts one test per direction. The edge ranking procedure performs $2m$ tests and fits $4m$ regression models, while the orientation procedure performs at most $|U|=m$ tests and fits $4m$ models. However, there are fewer tests and models required in practice because Meek’s rules will orient additional edges. The computational complexity of procedure OrientEdges is then of order $O(m)$. In the edge deletion step, the method performs one significance test per node on its parent nodes, amounting to $p$ tests and $p$ models, and has a complexity of order $O(p)$. While the PC algorithm only conducts conditional independence tests and is exponential with respect to $p$ in the worst case, it becomes polynomial when the underlying DAG is sparse. The empirical runtime comparisons are provided in Section 6.3.
6 Numerical Experiments
@@
We conducted numerical experiments with synthesized data to benchmark the accuracy and effectiveness of our method. At a detailed level, we assess the performance of the ranking procedure in Section 6.1, as well as the type I error rate and statistical power of the likelihood ratio test in Section 6.2. We then compare our method to competing causal discovery algorithms using simulated data sets in Sections 6.3. Intermediate results from each stage of our algorithm are provided in Section 6.4 to illustrate the effects of the individual components. Two real-world applications are presented in Section 7.
@@
We develop two variations of the likelihood ratio test in our algorithm: the sample-splitting (SNOE-SS) approach, which is delineated in Algorithm 4, and the cross-validation (SNOE-CV) approach. The CV approach employs two-fold cross-validation in Algorithm 4 to perform the likelihood ratio test twice by exchanging the training and test data sets and uses either the smaller or larger p-value for evaluation. The larger p-value is used in our experiments, but practitioners may specify either option. To learn the initial graph, we implemented our modified version of the PC-stable algorithm (pc_stable) from the bnlearn package coupled with the partial correlation test (scutari2010learning). A more stringent threshold of $\alpha_{1}=0.05$ was applied for learning the CPDAG, while a relaxed threshold of $\alpha_{2}=0.25$ was used for obtaining the additional candidate edges. The significance level for the likelihood ratio test was set at $\alpha=0.05$ and further reduced to $\alpha=10^{-4}$ for edge pruning. The algorithm is implemented as an R package and can be accessed at https://github.com/stehuang/snoe.git.
@@ 6.1 Accuracy of Ranking Procedure
We performed several experiments to verify the precision of our edge ranking procedure. Specifically, we tested its ability to correctly rank undirected edges in a CPDAG. Three distinct DAG structures were considered, as depicted in Figure 4, with $N=200$ data sets generated for each structure, each containing $n=2000$ samples. As the data was generated using nonlinear functions, the edge directions can be determined under all settings. Given the CPDAG of each DAG, we computed the edge-wise independence measure $\widetilde{I}(X,Y)$, defined in (8), for each undirected edge XYX-Y and ranked the edges in ascending order. We assessed whether an edge satisfying the PANM was ranked first.
AABBZZXXYYDAG 1
AABBZZXXYYDAG 2
AABBZZXXYYWWVVDAG 3
@@
Results are presented in Table 1 and are categorized by the data generating function used in the simulations. Note that the cubic, piecewise linear, and sigmoid functions are invertible. The values represent the proportion of data sets in which an undirected edge was ranked first for orientation. Uniformly across all graphs and functions, the edge following the PANM has the highest proportion of being ranked first. It is identified in the vast majority of data sets generated by the quadratic and sigmoid functions. Notably, our ranking procedure successfully determined the edge that satisfies the PANM in most cases under the cubic and piecewise functions, which is more challenging to distinguish since these invertible functions can be well-approximated by a linear function. These experiments verify both the ranking procedure and the use of the edge-wise independence measure to identify an edge fulfilling the orientation criterion.
@@
1 ZXZ-X Yes 0.64 0.62 0.99 0.72
1 XYX-Y No 0.36 0.38 0.01 0.28
2 ZXZ-X Yes 0.43 0.52 0.46 0.93
2 ZYZ-Y No 0.34 0.25 0.35 0.07
2 XYX-Y No 0.23 0.23 0.19 0
3 ZXZ-X Yes 0.38 0.52 0.95 0.41
3 XYX-Y No 0.14 0.24 0.01 0.19
3 YWY-W No 0.26 0.14 0.01 0.21
3 WVW-V No 0.22 0.11 0.03 0.19
Table 1: Frequency of an undirected edge ranked first under various settings.
@@ 6.2 Evaluation of the Likelihood Ratio Test
In this subsection, we investigate the type I error and statistical power of the likelihood ratio test. We considered five disinct DAG structures, each comprising of 2 to 5 nodes and 1 to 7 edges, and generated data sets of sample sizes $n=250,500,1000,1500,2000$. In the CPDAG of each network, we applied the likelihood ratio test to a targeted undirected edge to determine its causal direction. A total of $N=400$ tests were performed per graphical structure and sample size setting. Under a linear Gaussian DAG, the true edge direction of the targeted undirected edge is not identifiable. Under a nonlinear DAG, the true edge direction is identifiable.
@@
The type I error of the test, averaged across all DAG structures per sample size, is documented in Figure 5. Overall, the type I error deviates minimally from the specified significance level and stabilizes as the sample size grows. Under both significance levels, the maximal difference between $\alpha$ and the type I error is within 0.015 for both the sample-splitting and cross-validation approaches. This also signifies that the test is robust against type I errors at smaller sample sizes, which can be attributed to the sample splitting design that renders independence between the estimated model and test data. The difference between the two approaches is minimal as well, differing by 0.005 at $n=2000$, and indicates that both effectively control the false positive rate.
@@
A type II error under the likelihood ratio test would be to falsely declare the two models as equivalent when only one model is true. In regards to edge orientation, that is to fail at identifying the true causal direction of an edge or incorrectly orient an edge. For this experiment, we applied the likelihood ratio test to data generated from a singular non-linear function, where the orientation of the targeted edge is identifiable. We recorded the power of the test under significance level $\alpha=0.05$.
As seen in Figure 6, the statistical power of the likelihood ratio test increases with the sample size across all function types. For certain functions like the piecewise and secant functions, the power approaches one with $n\geq 1500$ samples. For other functions, the power improves more gradually but still strictly increases with the sample size, including the case of the cubic function. In particular, the power increases by at least 33% between sample sizes $n=250$ and $n=1000$. The cross-validation based approach exhibits greater power, achieving at least 85% by $n=2000$ across all function types. The results provide empirical evidence that the test can in practice identify the true causal direction of an undirected edge in a CPDAG, especially with sufficient data.
Figure 6: The statistical power of the likelihood ratio test on an undirected edge in a CPDAG, with select nonlinear functions underlying the SEM. Results show that power increases as $n$ increases to at least 80%.
@@ 6.3 Comparison of Algorithm Performances
We evaluated the performance of SNOE against competing methods on synthetic data. We compared our method to CAM, NOTEARS, DAGMA, and SCORE, where each represents a different approach for learning nonlinear DAGs. CAM (cam) is a score-based method that assumes a nonlinear causal additive model (4) with Gaussian noise and optimizes the log-likelihood function to learn a DAG. Utilizing deep neural networks to model SEMs, NOTEARS (zheng2020learning) and DAGMA (bello2023dagmalearningdagsmmatrices) formulate the structural learning problem as a continuous-optimization problem with an algebraic constraint to enforce acyclicity. In the following experiments, we employed the version tailored to learning from nonlinear data for both algorithms. SCORE (rolland2022score) employs a bottoms-up approach to iteratively identify leaf nodes by computing the Jacobian of the score function under the assumption of a Gaussian error distribution. For all methods, we used their recommended parameter settings. While several constraint-based methods were tested as well, their performance fell short. A detailed analysis of their performances is provided in Appendix C.1.
The algorithms were applied to learn six DAG structures of varying sizes selected from the bnlearn network repository. For each network, we generated $N=75$ data sets, each with a sample size of $n=1000$, using the additive model with Gaussian noise in Equation 4. The SEMs were created under three separate functional forms: linear functions, invertible functions, and non-invertible functions. The function classes are denoted respectively as linear, inv, and ninv in the figures. Under the invertible functions setting, we randomly selected functions from a set consisting of the cubic, inverse sine, piece-wise linear, and exponential functions. The non-invertible functions were sampled from a Gaussian process using a squared exponential kernel and bandwidth $h\sim\text{Unif}(5,5.25)$. For all cases, the Gaussian noise term was sampled with mean $\mu=0$ and standard deviation $\sigma\sim\text{Unif}(0.5,0.75)$.
The true DAG serves as the ground truth for evaluation, with the exception of the linear, Gaussian case for which only the MEC is identifiable and thus the true CPDAG is used as the ground truth. Our main evaluation metrics are the F1 score, structural Hamming distance (SHD), and computational complexity. The F1 score is a harmonic mean of the precision and recall scores. It is calculated as $F1=\frac{2TP}{2TP+FP+FN+IO}$, where TP, FP, FN, and IO respectively denote the number of true positives, false positives, false negatives, and incorrectly oriented edges. The SHD measures the number of edge additions, deletions, and reversals required to convert the learned DAG into the true DAG. The computational complexity is measured by overall runtime in seconds.
@@
The results presented in Figure 7 demonstrate that SNOE consistently outperforms competing methods. We observe that our algorithm achieved uniformly high F1 scores across all network structures and functional forms, with an average standard deviation of 0.06 in its performance across the three types of functions for both approaches. SNOE performed particularly well on invertible nonlinear DAGs, which presents a more challenging task due to the difficulty of detecting such nonlinear relations. On average across all function types, the F1 score of SNOE is respectively 67.6%, 61.8%, and 100.1% higher than those of NOTEARS, DAGMA, and SCORE. A closer analysis reveals that NOTEARS and DAGMA produced sparser DAGs by missing considerable amounts of edges, while SCORE often included many extraneous edges without capturing the true edges. CAM performed similarly to our method under data generated from non-invertible functions but showed considerable variability based on the data-generating function. Its F1 scores for invertible and linear functions are lower than SNOE with sizable margins for most networks. Specifically, CAM produced denser DAGs with more false positives, especially under these settings.
Our method also ran significantly faster than the competing methods. We show the average runtime on the $\log_{10}$ scale against network size (number of nodes) in Figure 8. The sample-splitting approach (SNOE-SS) has a slightly shorter runtime than the cross-validation approach (SNOE-CV), since the latter performs regression twice. For the largest network, our algorithm was at least 7.7 times faster than all competing methods. While CAM showed similar F1 scores in certain cases, SNOE-SS was between 2.8 to 10.7 times faster than CAM, with the difference magnified when learning larger networks. This efficiency can be attributed to its local learning approach of identifying edges satisfying the PANM, rather than optimizing over the entire DAG. Further reduction in runtime occurs as Meek’s rules typically orient additional undirected edges after performing the orientation test. In contrast to score-based and optimization-based methods that search over a restricted DAG space, our method leverages properties inherent to the graphical structure and only evaluates a sub-DAG, containing just the relevant nodes, to determine the correct edge orientations. Consequently, this results in higher computational efficiency for structural learning.
Figure 8: Average runtime in $\log_{10}(\text{seconds})$ of algorithms by network size.
@@
We further investigated the empirical performance of our algorithm under model mis-specification, specifically when the noise distribution is non-Gaussian, and present results in Figure 9. Recall that the DAG is identifiable when the noise terms follow a non-Gaussian distribution. For this experiment, we simulated data under the same previous settings, but sampled the error variables from three different non-Gaussian distributions: the t-distribution with $df=5$, the Laplace distribution, and the Gumbel distribution, all with $\mu=0$ and $\sigma\sim\text{Unif}(0.5,0.75)$. Since the learning accuracies for each error distribution are similar, we present the combined results. Both variations of SNOE achieved higher accuracy than competing methods across all settings again, while only CAM was comparable in a few cases. Similar to the Gaussian case, the F1 score of our method is consistent under different function types with non-Gaussian noise. The exact F1 scores are comparable to the Gaussian case as well, indicating that our method is robust to model mis-specification and is a versatile causal learning method.